OpenStudy (anonymous):

Y varies directly with x and y= -9 when x= -45

OpenStudy (igreen):

What are you trying to find? The constant?

OpenStudy (anonymous):

What is the direct variation equation

OpenStudy (anonymous):

I think it is. Y=1/5x

OpenStudy (igreen):

Direct variation equation: y = kx Where 'k' is the constant. Plug in what we know: -9 = k(-45) -9 = -45k Divide -45 to both sides: -9/-45 = k 1/5 = k So we have: y = 1/5x You are correct! \(\Huge\color{lime}\checkmark\)

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

Thank yoy